How many Pennies and nickels do u need to make $520in cash?
Answer by Theo(13342)

$520 is equal to 520 dollars.

each dollar contains 100 cents, so 520 dollars is equal to 52,000 cents.

each nickel is worth 5 cents, so take 52,000 and divide it by 5 and you get 10,400 nickels.

each penny is worth 1 cent, so take 52,000 and divide it by 1 and you get 52,000 pennies.

any combination in between will do as long as you work it by changing the values 5 cents at a time.

for example:

you can have 10,400 nickels and 0 pennies.
take away one nickel and you have 10399 nickels and 5 pennies.
take away another nickel and you have 10398 nickels and 10 pennies.

so the number of nickels and pennies can be any combination of nickels from 0 to 10,400 and any combination of pennies from 0 to 52,000.

you can actually graph this.

start with 5x + y = 52,000.

x is the number of nickels and y is the number of pennies.

solve for y to get y = 52,000 - 5x.

graph that equation.

you can see from the graph that,...

when x = 0, y = 52,000
when x = 6,000, y = 22,000
when x = 10,400, y = 0

x = 0 and y = 52,000 tells you that there are 0 nickels and 52,000 pennies.
since each nickel is worth 5 cents and each penny is worth 1 cent, the total value is 0 * 5 + 52,000 * 1 = 52,000 cents which is equal to 520 dollars

x = 6,000 and y = 22,000 tells you that there are 6,000 nickels and 22,000 pennies.
since each nickel is worth 5 cents and each penny is worth 1 cent, the total value is 6,000 * 5 + 22,000 * 1 = 52,000 cents which is equal to 520 dollars.

x = 10,400 and y = 0 tells you that there are 10,400 nickels and 0 pennies.
since each nickel is worth 5 cents and each penny is worth 1 cent, the total value is equal to 10,400 * 5 + 0 * 1 = 52,000 cents which is equal to 520 dollars.

you can pick any value of x and the graph will tell you the corresponding value of y.

you can also use the formula without graphing.

the same formula applies.

y = 52,000 - 5*x.

x tells you the number of nickels, which you supply between 0 and 10,400.
y tells you the number of pennies.

this is what the graph will look like:

$$$